OpenStudy (anonymous):

What are the coordinates of the midpoint of TD with T (-1, 5) and D ( 3, -9)?

OpenStudy (anonymous):

(1,-2)

OpenStudy (anonymous):

use the midpoint formula: (X1 + X2)/2 , (Y1 +Y2)/2 = ( -1 + 3) /2 , (5+ -9)/2 = (1, -2)